### v1

# Verb Mirage: Unveiling and Assessing Verb Concept Hallucinations in Multimodal Large Language Models

Abstract
Multimodal Large Language Models (MLLMs) have garnered significant attention recently and demonstrate outstanding capabilities in various tasks such as OCR, VQA, captioning, etc. However, hallucination remains a persistent issue. While numerous methods have been proposed to mitigate hallucinations, achieving notable improvements, these methods primarily focus on mitigating hallucinations about object/noun-related concepts. Verb concepts, crucial for understanding human actions, have been largely overlooked. In this paper, to the best of our knowledge, we are the first to investigate the verb hallucination phenomenon of MLLMs from various perspectives. Our findings reveal that most state-of-the-art MLLMs suffer from severe verb hallucination. To assess the effectiveness of existing mitigation methods for object concept hallucination on verb hallucination, we evaluated these methods and found that they do not effectively address verb hallucination. To address this issue, we propose a novel rich verb knowledge-based tuning method to mitigate verb hallucination. The experiment results demonstrate that our method significantly reduces hallucinations related to verbs. Our code and data will be made publicly available.
1 Introduction
Multimodal Large Language Models (MLLMs) [zhu2024minigpt, dai2023instructblip, liu2024improved, chen2024internvl, bai2023qwen, ye2024mplug] have drawn much attention in both research and industry community. Armed with high-quality data, a large number of parameters, and efficient instruction-following finetuning [liu2023visual], they achieve great success in many tasks such as OCR, VQA, and image captioning, demonstrating strong generalization ability.
Figure 1: Besides the well-discussed object hallucination, in this paper, we unveil the severe verb hallucination of state-of-the-art MLLMs with our designed benchmarks via multiple choice questions (MCQ) and binary questions (YN). Gemini-1.5-Flash and GPT-4-Turbo are tested with 100 randomly sampled questions.
However, MLLMs’ performance improvement could be hindered by hallucination. Typically, hallucination [ji2023survey, liu2024survey] means the output of MLLM contains contents against facts, irrelevant or nonsensical given context such as prompt or multimodal input. To test MLLM hallucination in different tasks, many benchmarks [liu2024mmbench, guan2024hallusionbench, cai2023benchlmm, fu2023mme] have been made, allowing people to assess MLLMs’ abilities in various aspects. To mitigate MLLM hallucination, many methods [huang2024opera, leng2024mitigating, sun2023aligning, yin2023woodpecker, zhou2023analyzing] have been proposed, successfully relieving hallucination to a large extent.
However, existing benchmarks and methods mainly target hallucination about objects/noun-related concepts. Action/verb-related concepts, which are crucial to understanding human actions, are typically overlooked. To this end, we propose to dig into the verb hallucination problem. We build the first verb-hallucination-oriented benchmark, which is based on existing datasets [chao2015hico, sigurdsson2018actor] needless of extra manual annotations. As MLLMs are a cooperation of vision and language modalities, we probe MLLM verb hallucination given both different visual inputs and language inputs, covering different query conditions, different imaging conditions, and different semantic conditions. Extensive experiments show that all MLLMs perform poorly on many aspects and thus show severe verb hallucination.
Moreover, we test existing low-cost hallucination mitigation methods on widely used MLLMs and show that they all fail in mitigating verb hallucination. To somewhat relieve verb hallucination, we propose a baseline method based on parameter-efficient fine-tuning with verb structure knowledge. Experiments show that our method successfully relieves verb hallucination, but its performance is still far from satisfactory. Finally, we explore the reason for verb hallucination and discuss possible future solutions.
In conclusion, our contributions are:
  1. 1.
    To our knowledge, we point out and analyze MLLM’s verb hallucination for the first time and probe this phenomenon from different aspects.
  2. 2.
    We evaluate existing low-cost hallucination mitigation methods on widely used MLLMs, and find out they all fail to alleviate verb hallucination.
  3. 3.
    We propose a baseline that can relieve verb hallucination better and discuss the reason for verb hallucination and possible solutions, paving the way for eliminating verb hallucination in future work.
2 Related Work
2.1 MLLM Benchmarks
Before the emergence of LLMs, great efforts have been made to build datasets focused on specific tasks like image captioning [chen2015microsoft, sharma2018conceptual, young2014image], VQA [goyal2017making, hudson2019gqa, marino2019ok], OCR [singh2019towards], action recognition and detection [tincvpr, chao2015hico, djrn, idn, sigurdsson2018actor, gu2018ava, pastanet], etc. However, they mainly assess the abilities of domain-specific tasks. To fully evaluate the abilities of MLLMs, more benchmarks have been proposed [fu2023mme, liu2024mmbench, li2024seed, ying2024mmtbench] to test different aspects and subtasks. Benchmarks are also proposed to conduct detailed assessments on specific aspects, like BenchLMM [cai2023benchlmm] for robustness against image styles, MMSpuBench [ye2024mm] for robustness against spurious correlations, and HaloQuest [wang2024haloquest] for MLLM multimodal reasoning capabilities.
2.2 MLLM Hallucination
Among the emerging benchmarks, hallucination has become a major focus. Typically, hallucination means that the contents generated by models are untruthful, against facts, or nonsensical [ji2023survey, liu2024survey, zhang2023siren]. Many benchmarks targeting hallucination have been proposed [guan2024hallusionbench, ye2024mm, wang2024haloquest]. In POPE [li2023evaluating], binary questions are used to probe hallucination about a certain class of object. CHAIR score [rohrbach2018object] is used to measure object hallucination of MLLMs in image captioning. Researchers also study other phenomena, such as event hallucination [jiang2024hal], hallucination snowballing [zhong2024investigating], and relation hallucination [wu2024evaluating]. The cause of hallucination is also investigated. OPERA [huang2024opera] recognizes MLLMs’ over-attention to summarizing text tokens as an outstanding phenomenon of object hallucination. VCD [leng2024mitigating] reveals that language prior is an important factor in hallucination. However, in most previous methods, only hallucinations involving objects were covered by identifying whether the output of MLLM contains a reference to an object that does not match the image or appears in the image. Verb-related concepts, which are crucial to understanding human activities, were neglected. Instead, we omit relevance to object hallucination, focus on verb concepts, and probe verb hallucination on different vision and language conditions.
2.3 Hallucination Mitigation
Researchers have revealed the reason for hallucination from many different aspects and proposed hallucination mitigation methods respectively. To omit the bias or errors in training data, researchers proposed mitigating bias [liu2024mitigating, hu2023ciem, gunjal2024detecting] in the dataset or enriching the annotation [zhai2023halle]. Some works mitigated hallucination by scaling up the resolution of input images [bai2023qwen, chen2024internvl, li2023monkey]. Moreover, some suggest post-processing at inference time by adjusting decoding strategy [huang2024opera, leng2024mitigating, chen2424halc] or correcting the output of MLLMs with the help of expert models [zhou2023analyzing, yin2023woodpecker].
3 Probing on MLLM IO Conditions
We probe verb hallucination from different perspectives, e.g., MLLM behavior given different question formats, image qualities, verb semantics, and angles of view, etc.
We select HICO [chao2015hico] and CharadesEgo [sigurdsson2018actor] as the main datasets for probing MLLM verb hallucination. HICO contains 47K images with dense annotations. It includes 600 action classes formed by 80 object classes and 117 verb classes. It has rich verb labels and is thus suitable for evaluating MLLM verb understanding with minor manual adaptation. Meanwhile, CharadesEgo contains 7K videos of daily indoor activities. In each scenario, the same actor is recorded with both an egocentric and exocentric camera. It contains 157 action classes, each formed by a verb and an object.
We test several open-sourced and close-sourced MLLMs including InstructBLIP-7B [liu2023visual], LLaVA-V1.5-7B [liu2024improved], mPLUG-Owl2 [ye2024mplug], Qwen-VL-Chat [bai2023qwen], MiniCPM-Llama3-V2.5 [yao2024minicpm], Qwen2-VL-7B-Instruct [Qwen2VL], GPT-4-Turbo [achiam2023gpt], Gemini-1.5-Flash [team2023gemini], etc. They have different ranks on commonly used leaderboards and show outstanding results on benchmarks targeting object concepts.
Figure 2: We probe MLLM verb hallucination from various perspectives, e.g., question formats, the existence of object correlation, different fields of view, image qualities, verb semantics, and image semantics.
3.1 Probing on Different Query Conditions
3.1.1 Different Question Formats
Nearly all MLLMs follow the VQA format. Therefore, users can interact with MLLMs naturally by asking questions and requiring answers from MLLMs in natural language. MLLMs with low hallucinations should give hallucination-free answers given different question formats. Therefore, we evaluate MLLM hallucination using different question formats including Multiple Choice (MC) questions with only one correct answer each and Yes-or-No (YN) questions. Here we do not introduce free-form image captioning and blank-filling because these two forms require rule-based post-processing on verbs and may lead to severe misclassification errors.
When building MC questions, for a sample image, we randomly choose a verb presented in the image and three verbs possibly performable upon objects but not presented. We introduce circular evaluation [liu2024mmbench] and regard accuracy as a metric for MLLM verb hallucination. For YN questions, when building negative samples, we also randomly choose possible verbs for the objects but not carried out in the image. We regard accuracy, precision, recall, and F1 score as the metrics for hallucination.
We aim to verify none other than verb hallucination, so we do our best to omit relevance to object hallucination. For each verb-object tuple, we form questions by altering the verb and leaving the object unchanged. For example, if an image contains a person holding a cup, we may ask MLLM, “Is someone holding a cup? Is someone washing a cup?” In this way, relevance with object hallucination can be minimized. However, we must point out that as a substantial proportion of verbs are transitive verbs, the influence of objects can not be completely omitted.
3.1.2 Object Correlation
Sometimes we focus on human interaction with a certain class of object, but sometimes our focus on verbs may be object class agnostic. Specifically, we may wonder “Is someone holding a cup in the image?” However, sometimes we may also want to know “Is someone eating something in the image?” Therefore, we test MLLM verb understanding both given reference to objects and not. Among these two conditions, we believe questions without object correlation (i.e., “Is someone eating something?”) have less relevance to object hallucination, but questions with object correlation are also very practicable in daily use.
3.1.3 Analysis and Discussion
The results are shown in Tab. 1, giving us rich clues about MLLM verb hallucination.
Heavy reliance on objects. MLLMs show drastic performance degradation on MC questions without reference to objects. Detailed statistics on YN questions based on object classes referred to in the questions also reveal that MLLM verb understanding relies heavily on object reference. We analyze some commonly used datasets [sharma2018conceptual, changpinyo2021conceptual, lin2014microsoft, ordonez2011im2text, krishna2017visual, hudson2019gqa, goyal2017making] for MLLM pretraining and investigate the number of nouns and verbs in the datasets in Fig. 3(a). We can see that the number of nouns is 4-10 times the number of verbs. One reason for this unbalanced ratio of nouns and verbs is that datasets on action understanding have not attracted enough attention from the MLLM community. Another reason is that there are many more nouns than verbs in English. Research on shortcut learning [geirhos2020shortcut] also sheds light on the reason for MLLMs’ overreliance on nouns.
Inability to refuse. All MLLMs have high recall but low precision, meaning that MLLMs tend to give Yes whether a certain verb is presented in the image or not. Binary questions require MLLMs to have a deep understanding of verb concepts in images, which are more difficult to answer than MC questions, but vitally important in daily use.
YN w/ obj YN verb only MC w/ obj MC verb only
Model acc prec recall acc prec recall acc acc
Qwen2-VL-7B 75.51 58.37 93.75 74.69 57.43 95.87 71.47 65.31
MiniCPM-Llama3-V2.5 80.91 66.83 85.41 79.14 63.33 90.33 66.39 60.77
Qwen-VL-Chat 78.06 62.37 87.02 79.24 65.09 82.68 55.95 54.57
mPLUG-Owl-2 62.94 47.38 95.99 62.61 47.25 94.94 63.91 62.60
LLaVA V1.5 49.50 39.80 98.94 58.21 44.49 97.49 57.37 51.00
InstructBLIP 72.53 55.79 86.77 73.82 57.25 87.79 13.48 6.25
Table 1: Results on YN and MC questions w/ and w/o object reference. Red: high recall. Blue: low precision. Bold: higher MC acc w/ object reference than w/o object referece.
Figure 3: Comparison between objects and verbs.
3.2 Probing on Different Imaging Conditions
3.2.1 High-Quality and Low-Quality Images
Previous research [li2022hake, leng2024mitigating] has revealed that visually distorted images can hinder both humans and models from recognizing the contents in images well. However, the relation between visual distortion and verb hallucination is unexplored. Do MLLMs hallucinate in the same way when given high-quality images and visually distorted images? Is verb understanding more sensitive to visual distortion than object understanding for MLLMs? Here, we add pepper-salt noise as visual distortion to images for testing.
We evaluate MLLM verb understanding with both high-quality and visually-distorted images and report performance and error consistency following [geirhos2021partial] in Tab. 2. All tested MLLMs show obvious performance degradation. Error consistency in the form of Cohen’s Kappa [cohen1960coefficient] measures MLLM consistency of answers given different visual conditions and provides a guideline for MLLM performance improvement. We can see some MLLMs with low ranks do not have bad error consistency, but MLLMs with higher ranks show low error consistency. This means that their verb hallucination can be easily induced by visual distortion.
As a control experiment, besides verb understanding, we also build a test set for MLLMs’ object understanding with the same set of images. The observation consistency in terms of Cohen’s Kappa is reported in Fig. 3(b). From the result, we can see that MLLM shows much higher inconsistency in verb understanding than object understanding, which means that visual distortion does more harm to verb understanding than object understanding.
In conclusion, we have the following findings:
Finding: Visual distortion affects both object understanding and verb understanding of MLLMs, but the effect on verb understanding is greater.
YN verb only MC verb only
w/o Pepper Salt w/ Pepper Salt YN Err. Cons. w/o Pepper Salt w/ Pepper Salt MC Err. Cons.
YN acc YN prec YN recall YN acc YN prec YN recall MC acc MC acc
Qwen2-VL-7B 74.69 57.43 95.87 63.20 47.68 95.57 56.86 65.31 51.94 48.21
MiniCPM-Llama3-V2.5 79.14 63.33 90.33 67.40 51.25 64.79 26.12 60.77 40.50 37.20
Qwen-VL-Chat 79.24 65.09 82.68 66.64 50.28 80.43 38.47 54.57 33.98 43.38
LLaVA V1.5 58.21 44.49 97.49 51.29 40.66 97.30 73.85 51.00 49.97 68.37
InstructBLIP 73.82 57.25 87.79 71.04 54.35 87.40 74.16 6.25 6.34 82.33
Table 2: Performance comparison for images w/ and w/o pepper-salt noise. Underline: higher performance w/o peppser-salt noise. Red/Blue: good/bad error consistency (Err. Cons.).
3.2.2 Egocentic and Exocentric Images
Recently, MLLMs have shown outstanding capabilities in recognition and reasoning, and a growing body of research has delved into leveraging MLLMs for various tasks in Robotics, where egocentric images are widely used. Therefore, a study of MLLM’s understanding of verbs in egocentric images holds significant importance. To evaluate MLLMs’ understanding of verbs in egocentric images and the performance gap between egocentric and exocentric images, we build a small test set using Charades-Ego and conduct experiments on different MLLMs.
Given the same scenario recorded with exocentric and egocentric cameras, we also probe MLLMs’ understanding with MC and YN questions. The results are shown in Fig. 4 and Tab. 3, there is a substantial gap between MLLMs’ understanding of exocentric and egocentric images.
Finding: MLLMs can not understand verb concepts in egocentric images as well as exocentric images. MLLMs are better at exocentric verb understanding.
Figure 4: Performance comparison on egocentric and exocentric verb understanding (question type: MCQ).
View Egocentric Exocentric
# of Samples 13.6K 14.5K
Model acc prec recall acc prec recall
Qwen2-VL-7B 60.10 60.31 59.12 63.01 60.98 72.25
MiniCPM-Llama3-V2.5 59.42 62.60 46.78 62.00 67.44 46.39
Qwen-VL-Chat 57.06 61.87 36.79 60.62 64.94 46.19
LLaVA V1.5 56.29 54.50 76.28 58.91 56.21 80.60
InstructBLIP 57.59 60.45 43.87 59.86 62.11 50.55
Table 3: Comparison on egocentric and exocentric verb understanding (question type: YN).
3.3 Probing on Different Semantic Conditions
3.3.1 Rare and Common Verbs
Verbs follow long-tailed distribution in action datasets because of all sorts of difficulties in the process of dataset collection. However, understanding rare and common verbs is equally important in real-world applications. We hypothesize that MLLMs tend to hallucinate more on rare verbs and try to prove it on existing datasets.
We divide the negative samples into two subsets: the rare set and the common set. In rare set, the verb in question lies in the tail of HICO verb distribution while in common set, the verb in question lies in the head. Specifically, for YN questions with an object reference, the rare set contains all HOIs whose annotations make up less than 20% among HOIs relevant to the same object class. For YN questions without object reference, the common set contains all questions containing hold, ride, sit_on, straddle and carry, making up 50% of the verb annotations in HICO dataset. From Fig. 5, we can see that MLLMs tend to refuse existent rare verbs but accept nonexistent common verbs in images. This phenomenon reveals that the long-tailed distribution of verb annotations limits MLLM verb understanding. How to understand rare verbs remains a problem and there is a large room for action data collection and curation.
Finding: MLLMs can not understand rare verbs as well as common verbs, i.e., long-tail affects a lot.
Figure 5: True negative correct rate comparison on rare and common subsets.
3.3.2 Image Content Ambiguity
Ambiguity always exists in real-world scenarios. Understanding verbs in crowded or heavily occluded scenarios is important in many fields such as surveillance, social robotics, and visual reasoning. To assess MLLMs’ verb understanding given images with ambiguous content, we select images from HICO containing content ambiguities, form an ambiguous subset, and compare them with images with less content ambiguity. Specifically, our ambiguous subset contains many contributing factors to ambiguity:
  • Imbalanced human-object relative size. Imbalanced human-object relative sizes can add difficulties to MLLM perception of humans and objects. The existence of verbs relies heavily on the accurate perception of humans and objects in images. Potential failures in perception bring great difficulties to the recognition of verbs.
  • Crowded scene. A highly complicated scene structure can distract MLLMs. To judge the existence of a verb, MLLMs must analyze all humans and objects and get a comprehensive conclusion. The large number of humans and objects puts heavy burdens on MLLMs.
  • Occlusion. We can recognize humans and objects according to part of them and analyze their relationship even with prominent occlusion. Thus, visual reasoning under occluded scenarios is important in fields like action understanding and robot manipulation.
From Fig. 6, we have the following finding:
Finding: MLLMs show performance degradation when images contain content ambiguity. How to understand verbs given ambiguous content is an open problem to be solved.
Figure 6: MLLM accuracy on ambiguous (Amb.) and unambiguous (Unamb.) subsets.
4 Probing on Model Behaviors
In this section, we probe verb hallucination from the model perspective.
4.1 Token Uncertainty
We first dig into the uncertainty [zhou2023analyzing] of MLLMs via visualizing the distribution of probabilities of predicted tokens of the widely used open-sourced LLaVA V1.5. From Fig. 7, we can see that there is a substantial difference between correct answers and hallucinated answers: hallucinated tokens are mostly given with low probability, which means that the model is confused about the answers. For Yes-or-No questions, we visualize questions to which the model gives answers Yes and No separately. The observations are:
  1. 1.
    LLaVA V1.5 tends to answer Yes with high confidence, but No with relatively lower confidence.
  2. 2.
    Non-hallucinated answers are given with higher confidence than hallucinated ones, regardless of Yes and No.
  3. 3.
    For MC questions, answers given by LLaVA V1.5 with higher confidence are very likely to be correct.
Finding: MLLM verb hallucination has something in common with MLLM object hallucination. Uncertainty is strongly related to MLLM hallucination.
Figure 7: Probability distribution of hallucinated tokens and not hallucinated tokens.
4.2 Attention
4.2.1 Key Image Area Attention
We hypothesize that models pay less attention to key information in images and text when they give hallucinated answers. We still take LLaVA V1.5 as an example and study the relationship between LLaVA V1.5 verb hallucination and its attention to tokens related to key areas in images. From Fig. 8(a), we can see that there is an obvious distinction of distribution between hallucinated attention and non-hallucinated attention, showing a strong correlation between inadequate attention to key areas and hallucination.
4.2.2 Verb Token Attention
Previous experiments give us a strong impression that MLLM relies too much on objects. Therefore, we hypothesize that MLLM tends to pay more attention to object tokens. In Fig. 8(b), MLLM’s attention to verb tokens is normalized by the sum of verb attention and object attention. As MLLM tends to answer Yes, we focus on questions with correct answers No. For negative YN questions with object references, hallucinated tokens are given with inadequate attention to the verb in question. In Fig. 8(c), verb and noun attention are normalized by their sum. MLLM tends to give more attention to object tokens than verb tokens.
Finding: MLLM tends to pay more attention to object tokens than verb tokens.
Figure 8: Attention for YN questions with object references.
4.3 MLLM Error Consistency
Consistency on MC and YN questions. From Tab. 1, we find that there is inconsistency in MLLMs’ performance on multiple-choice and yes-or-no tasks. However, it is an interesting question whether there is some consistency. Thus, we divide the questions into different classes according to object class references and get the object classes that influence MLLMs the most. Assume $S_{1}$ denotes the set of object classes without which MLLMs can not recall correct verbs given YN questions, and $S_{2}$ denotes the similar set of object classes for MC questions. We define overlap index=$\frac{|S_{1}\cap S_{2}|}{\text{min}\{|S_{1}|,|S_{2}|\}}$, and compute MLLMs’ overlap index. From Tab. 4, we find substantial overlap between $S_{1}$ and $S_{2}$. A substantial number of object classes can affect MLLM verb understanding of both question types.
Qwen-VL-Chat mPLUG-Owl2 LLaVA V1.5 InstructBLIP
overlap 95.35 79.17 82.14 100
Table 4: Influential object set overlap of MLLMs.
Consistency among models. We analyze models and get verb classes that are influenced by object references most according to per verb class Yes ratio. From Tab. 5, we can see that there is substantial overlap among different models. We also ensemble the answers given by models and show results in Fig. 1. There is no substantial improvement over the individual models, showing that the models share similar biases.
Finding: A substantial number of objects affect MLLM verb understanding on both question formats and all models. MLLMs show error consistency to some extent.
IoU Qwen-VL-Chat LLaVA V1.5 mPLUG-Owl2 InstructBLIP
Qwen-VL-Chat - 59.30 42.86 52.78
LLaVA V1.5 59.30 - 46.91 53.57
mPLUG-Owl2 42.86 46.91 - 41.54
InstructBLIP 52.78 53.57 41.54 -
Table 5: YN influenced verb set IOU.
5 Hallucination Mitigation Methods
5.1 Training-Free Methods
OPERA [huang2024opera], VCD [leng2024mitigating], and HALC [chen2424halc] are outstanding hallucination mitigation methods working during inference. They do not require finetuning, so they are low-cost and have more general applicability. We evaluate them on benchmarks formed by MC and YN questions.
We present the results of OPERA in Tab. 6, showing that OPERA does not always show improvement. Even when it shows improvement, the gain is marginal. Though it has been claimed that OPERA tries to punish overreliance on summary tokens and force more attention on visual tokens [huang2024opera], from Fig. 9 we find that for questions with correct answer No hallucinated models tend to give more attention to visual tokens. If the reward is given for attention to visual tokens, hallucination will be worsened. As we can not know the correct answer ahead of time, we can not decide whether to reward high attention to visual tokens.
Figure 9: Probability distribution of visual token attention for positive samples and negative samples.
VCD regards language prior as a hallucination-inducing factor, uses visual distortion to trigger MLLM’s reliance on language prior, and proposes contrastive decoding to eliminate hallucination. We present the results of VCD in Tab. 6. VCD shows an inconsistent effect on three models and thus fails on our benchmarks. To dig deeper, we compute Qwen-VL’s KL divergence between the original vocabulary distribution and contrasted vocabulary distribution $\text{KL}(p_{\theta}(y_{t}|v,x,y_{<t})||p_{vcd}(y_{t}|v,v^{\prime},x,y_{<t})))$ of 20, 000 samples, and find that the KL divergence of 18, 624 samples is 0, meaning that in our benchmarks there is no strong element of language prior which usually hides but can be triggered by visual distortion.
HALC argues that field of view is a strong factor affecting MLLM hallucination, and proposes contrastive decoding given different fields of view. As HALC is very time-consuming, we evaluate part of our benchmarks and present the results of HALC in Tab.  7. It also fails to show substantial improvement. On one hand, HALC focuses too much on objects. In our benchmarks, to minimize relevance to object hallucination, all object references are correct which limits the use of HALC. On the other hand, HALC’s failure shows that fields of view do not have much influence on MLLM verb understanding.
w/ obj w/o obj
Qwen-VL-Chat LLaVA V1.5 InstructBLIP Qwen-VL-Chat LLaVA V1.5 InstructBLIP
YN acc vanilla 78.06 49.5 72.53 79.24 58.21 73.82
w/ OPERA 78.03 42.46 73.08 79.19 54.45 74.01
w/ VCD 77.23 52.38 73.88 78.88 57.86 74.39
YN prec vanilla 62.37 39.8 55.79 65.09 44.49 57.25
w/ OPERA 62.32 36.76 56.38 64.98 42.38 57.45
w/ VCD 61.42 41.18 57.42 65.48 44.27 58.28
YN recall vanilla 87.02 98.94 86.77 82.68 97.49 87.79
w/ OPERA 87.07 99.54 86.85 82.81 98.33 87.62
w/ VCD 86.11 98.3 85.3 78.83 97.26 84.28
MC acc vanilla 55.95 57.37 13.48 54.57 51 6.25
w/ OPERA 55.63 57.28 13.96 54.17 51.13 6.28
w/ VCD 55.95 54.26 5.77 52.47 48.94 3.74
Table 6: Comparison on benchmarks w/ and w/o OPERA and VCD.
YoN verb only (20K samples) MC verb only (20K samples)
w/o HALC w/ HALC w/o HALC w/ HALC
acc prec recall acc prec recall acc acc
mPLUG-Owl-2 55.85 42.47 97.74 56.07 42.59 97.76 60.18 60.18
LLaVA V1.5 54.04 41.58 99.25 54.00 41.56 99.25 51.26 51.26
InstructBLIP 72.44 55.05 86.78 72.44 55.05 86.78 6.54 6.54
YoN w/ obj (20K samples) MC w/ obj (20K samples)
w/o HALC w/ HALC w/o HALC w/ HALC
acc prec recall acc prec recall acc acc
mPLUG-Owl2 62.53 46.54 95.99 62.64 46.61 95.86 61.78 61.78
LLaVA V1.5 52.04 40.53 99.05 52.03 40.53 99.05 56.92 56.92
InstructBLIP 73.98 57.12 82.88 73.98 57.12 82.88 18.26 18.26
Table 7: Comparison on benchmarks w/ or w/o HALC.
5.2 Mitigation via Finetuning
The above experiments reveal severe MLLM verb hallucination and show very different properties from noun hallucination: there is no strong language prior, no strong reliance on visual tokens, and no strong dependence on fields of view. These phenomena give us a strong impression that MLLM did not learn verbs well. To mitigate verb hallucination to some extent without sacrificing MLLM’s ability in other perspectives, we further explore verb hallucination mitigation via direct finetuning and propose a method based on verb structure knowledge.
We try to advance the MLLM with Pangea [li2024isolated], which organizes existing heterogeneous action datasets in a unified way. It builds a mapping from action labels to abstract verb semantics. 290 frequent verb nodes in VerbNet [verbnet] are selected and a one-to-290 mapping is built. It gives us a whole picture of diverse verbs and carries the structure knowledge of verb relationships. We select 60,000 samples from Pangea according to the proportion of the source dataset and build an instruction-tuning dataset. It contains 280 out of 290 nodes in Pangea P2S mapping and covers a wide range of verb semantics. To improve MLLM’s ability on verb understanding without sacrificing MLLM’s ability on other perspectives, we choose to finetune LLaVA V1.5 with LoRA [hu2022lora]. The results are shown in Tab. 8.
From the result, we can see outstanding improvement in MC questions. Although Pangea only contains rough action labels, it proves helpful to verb hallucination mitigation. In the future, mining more action data according to the structured verb semantics and finetuning MLLMs on them can be a promising way to mitigate verb hallucination.
Finding: Finetuning model with data roughly annotated but containing rich verb semantics proves helpful to MLLM verb hallucination.
not Finetuned Finetuned not Finetuned Finetuned
YN YN MC MC
acc prec recall acc prec recall acc acc
w/ obj 54.75 42.37 97.31 48.24 39.21 98.99 52.60 62.20
verb only 75.49 59.64 84.14 61.00 46.23 97.11 50.42 62.17
Table 8: Finetuning result of LLaVA V1.5.
6 Discussion
We mainly probe verb hallucination via MC and YN questions and leave free-form captioning out, because our experiments show that MLLMs tend to give as few verbs as possible. Thus, evaluating verb hallucination via metrics similar to CHAIR score [rohrbach2018object] is impractical: unmentioned existing objects will not cause punishment. Only mentioned non-existing objects cause punishment. Metrics like the CHAIR score can not provide a full perspective for verb hallucination. Another issue is the synonym of verb/action and hierarchy processing. The same action can be described in many different ways(e.g., doing housework and mopping floor), and establishing such synsets is still a problem. Based on these considerations, we do not choose free-form captioning as verb hallucination probation.
Given the above experiments, verb hallucination is quite different from object hallucination. Although MLLMs have shown some understanding of verbs, their performance is still far from satisfactory. Existing training-free hallucination mitigation methods show inconsistent results and can be said to fail on our benchmarks. It is still an open question what factor inside MLLMs can trigger MLLM verb hallucination. In our experiments, verb understanding does not always rely on visual tokens, and there is no strong language prior. Given these observations, we hypothesize that existing widely used MLLMs overfit on objects but underfit on verbs. In the future, study about MLLM visual representation can give us concrete proof.
Mitigating verb hallucination via finetuning is a more promising way. As the data collection is labor-intensive and finetuning consumes high resources, how to fully use existing data is vital but still a problem to be explored. Pangea contains rich verb semantics and provides a guideline for data collection and selection. In the future, a full use of Pangea P2S mapping can provide guidance. As existing datasets also contain annotations about verbs, maybe we can delicately reuse data so that we can mitigate verb hallucination without introducing a large amount of data.
7 Conclusion
In this paper, to the best of our knowledge, we first reveal MLLM verb hallucination and build a benchmark to probe it from various perspectives. Our experiment reveals that MLLMs suffer from severe verb hallucination in many ways, and existing training-free hallucination mitigation methods fail. The experiment shows that MLLM verb hallucination is quite different from object hallucination. Thus, we propose to finetune MLLM with LoRA using data of low quality but with rich verb semantics to mitigate verb hallucination. However, how to finetune existing models efficiently is still a problem to be explored. Moreover, whether there are more effective training-free verb hallucination mitigation methods is still an open problem.
Supplementary Material for Verb Mirage: Unveiling and Assessing Verb Concept Hallucinations in Multimodal Large Language Models
Appendix A Details on Benchmark Construction
A.1 Benchmark Based on HICO
HICO [chao2015hico] test set contains 9.8K images. We use the HICO test set for benchmark construction. It includes 600 action classes formed by 80 object classes and 117 verb classes. There is a special verb class “no_interaction”, and we leave out annotations about it.
Assume $\mathcal{H}$ denotes the set of 600 action classes, $\mathcal{V}$ denotes the set of verb classes, and $\mathcal{O}$ denotes the set of object classes. Each action class $h\in\mathcal{H}$ can be denoted as $(v,o)$ where $v\in V$ and $o\in\mathcal{O}$.
For MC question construction, given an image $I$ and its HOI annotation $\mathcal{H}_{I}$, we select $h_{0}=(v_{0},o_{0})\in\mathcal{H}_{I}$ as a positive option, select 3 negative options from the set of plausible but not presented action classes $\{(v,o)|o=o_{0}\land(v,o)\notin\mathcal{H}_{I}\}$, and form a MC question with one correct option. In our benchmark, there are 16.6K unique MC questions.
For YN question construction, we build questions with correct answers Yes and No similarly to MC question construction. In our benchmark, there are 47K unique YN questions. The ratio of questions with correct answers Yes and No is approximately 1:2.
A.2 Benchmark Based on CharadesEgo
CharadesEgo [sigurdsson2018actor] contains 157 action classes, each of which is formed by a verb and an object. However, many object classes in CharadesEgo can not match more than 2 verbs. Therefore, to build negative options, we sample from the set $\{(v,o)|o\in\mathcal{O}_{I}\land(v,o)\notin\mathcal{H}_{I}\}$, where $\mathcal{O}_{I}=\{o|(v,o)\in\mathcal{H}_{I}\}$. The ratio of questions with correct answers Yes and No is approximately 1:1. There are 76K unique YN questions and 36K unique MC questions, and we use 28K YN questions and 10K MC questions for experiment in the main paper.
Appendix B Additional Explanation and Analysis on Model Behaviors
B.1 Uncertainty
In the main paper, we visualize token uncertainty:
$\text{max}\{\text{softmax}(\text{logit}(y_{0}|v,x))\}$ (1)
of LLaVA V1.5 [liu2024improved]. In our setting, $y_{0}$’s are usually option tokens or Yes/No tokens, which are used to judge the correctness of MLLMs’ answers. Here we provide some additional visualizations in Fig. 10. The visualizations give us more solid proof that uncertainty is strongly related to MLLM verb hallucination.
Figure 10: Token uncertainty of InstructBLIP [dai2023instructblip]. Not hallucinated tokens are often given with high probability.
B.2 Key Image Area Attention
LLaVA V1.5 adopts CLIP [radford2021learning] as the visual encoder. 576 tokens generated by CLIP are used as visual tokens and fed to the language model after projection. Each of these 576 tokens represents a small area of the original image. Using the bounding box annotation of HICO, we can get the key image area of a certain verb, represented by the area bounded by the bounding boxes of humans and objects. Our visualization shows that hallucinated MLLM pays less attention to the key image area, but the margin is not large enough. Although MLLM can pay attention to the correct key areas, it can not often understand verb concepts.
B.3 Verb Token Attention
Verb Token Attention means the attention given to the verb tokens in question. When visualizing verb token attention, to emphasize LLaVA V1.5’s over-attention on noun tokens, we visualize verb token attention normalized by the sum of verb token attention and object token attention:
$\mathop{\text{mean}}_{j}\frac{\sum_{i\in T_{v}}\alpha_{ij}}{\sum_{i\in T_{v}}\alpha_{ij}+\sum_{i\in T_{o}}\alpha_{ij}},$ (2)
where $\alpha_{ij}$ represents the attention weight assigned to token $i$ at head $j$ in the last transformer layer, $T_{v}$ denotes the set of verb tokens, and $T_{o}$ denotes the set of object tokens.
B.4 Visual Token Attention
Visual Token Attention(VTA) is defined similarly to Visual Modality Contribution(VMC) [chen2024multi]. The Visual Token Attention is defined as
$\mathop{\text{mean}}_{j}\frac{\sum_{i\in V}\alpha_{ij}}{\sum_{i\in V}\alpha_{ij}+\sum_{k\in T}\alpha_{kj}},$ (3)
where $\alpha_{ij}$ represents the attention weight assigned to token $i$ at head $j$ in the last transformer layer, $V$ denotes the set of visual tokens, and $T$ denotes the set of textual tokens. In the implementation of OPERA [huang2024opera], when generating the first 10 tokens, a reward is given for high visual token attention. We visualize VTA on other models and other question types in Fig. 11 and  12, and have the following observations:
  1. 1.
    For questions with correct answers Yes, hallucinated models tend to give less attention to visual tokens.
  2. 2.
    For questions with correct answers No, hallucinated models tend to give more attention to visual tokens.
  3. 3.
    For MC questions, there is no substantial difference in visual token attention.
The observation above can explain the reason for the marginal or inconsistent effect of OPERA on MLLMs.
Figure 11: Visual token attention distribution for LLaVA V1.5.
Figure 12: Visual token attention distribution for InstructBLIP.
B.5 Object Reliance
We split YN question sets according to object classes, and show the difference between mPLUG-Owl2’s [ye2024mplug] accuracy with and without object references in Fig. 13. We can see that for questions about most object classes, mPLUG-Owl2 behaves differently with and without object references and thus is affected by object references although the accuracy does not change much (62.94/62.61).
Figure 13: mPLUG-Owl2’s difference of YN accuracy with and without object references.
Appendix C Details about Experiments on Training-free Hallucination Mitigation
C.1 Experiment Setting
Previous research [leng2024mitigating, chen2424halc] has demonstrated that sampling strategy can affect MLLM performance. Therefore, to evaluate the effect of hallucination mitigation methods fairly, we choose to control the use of decoding strategy. As OPERA [huang2024opera] heavily relies on beam search, we select beam search as the default decoding strategy for evaluation on OPERA and VCD [leng2024mitigating] in the main paper.
As HALC [chen2424halc] is very time-consuming and selects DoLA [chuang2024dola] beam search as the default sampling strategy in their codebase, we select DoLA beam search with beam size 1 as the default sampling strategy in our experiment.
C.2 Additional Experiments on VCD: Sampling as Decoding Strategy
We test VCD with sampling as the decoding strategy as the author claimed in the paper [leng2024mitigating], and get results in Tab.  9. From the result, we can see that VCD also shows inconsistent results. Sometimes sampling+VCD may achieve substantial improvement oversampling only, but the result is not better than beam search only.
w/ obj w/o obj
Qwen-VL-Chat LLaVA V1.5 InstructBLIP Qwen-VL-Chat LLaVA V1.5 InstructBLIP
YN acc sampling 77.55 51.46 65.96 78.94 52.21 65.34
sampling+VCD 76.68 52.14 68.59 78.79 53.12 68.67
beam search 78.06 49.50 72.53 79.24 58.21 73.82
YN prec sampling 61.63 40.56 49.5 64.49 41.09 49.08
sampling+VCD 60.72 40.98 52.02 65.34 41.64 52.28
beam search 62.37 39.8 55.79 65.09 44.49 57.25
YN recall sampling 87.4 96.37 78.65 83.29 96.73 77.98
sampling+VCD 86.06 97.35 80.52 78.79 97.74 79.59
beam search 87.02 98.94 86.77 82.68 97.49 87.79
MC acc sampling 55.74 37.15 0.9 54.26 39.29 0.73
sampling+VCD 55.74 46.41 1.55 52.09 47.47 1.01
beam search 55.95 57.37 13.48 54.57 51.00 6.25
Table 9: Comparison on sampling, sampling+VCD, and beam search. Bold: best among three. Underline: second best among three.
Appendix D Failure Case Visualization
In this section, we showcase some samples of which MLLMs hallucinate. MLLMs’ responses in red denote hallucinations. We can see MLLMs hallucinate because of inability to detect humans (Fig. 14(c) and Fig. 16(b)), inability to understand objects (Fig. 14(b), Fig. 15(d), and Fig. 16(d)), and inability to distinguish verbs (Fig. 14(a)(d), Fig. 15(a)(b)(c), and Fig. 16(a)(c)). A detailed analysis of MLLMs’ chains of thought can reveal that MLLMs can sometimes generate self-contradicting responses (Fig. 16(c)), imagine too much (Fig 14(d)), or improperly generalize concepts (Fig. 15(b)), which all lead to MLLM verb hallucination.
Figure 14: Some failure cases of MiniCPM-Llama3-V2.5 [yao2024minicpm].
Figure 15: Some failure cases of Gemini-1.5-Flash [team2023gemini].
Figure 16: Some failure cases of GPT-4-Turbo [achiam2023gpt].